No of oxalic acid molecules in  100 ml of 0.02 N oxalic  acid is 

Grade:11

2 Answers

Arun
4 years ago

Normality=molarity×basicity for acids/acidity for bases

Here N=0.02 & as oxalic acid is a dibasic acid so it will have a badicity of 2.

M=0.02÷2=0.01

Molarity=no of moles × volume(lit)

M=0.01×0.1=0.001moles

As 1 mole= 6.603×10*23=6.6×10*23×0.001= 6.6×10*20

Khimraj
4 years ago
 
As Oxalic acid is dibasic, n-factor (Basicity) = 2
Now, Normality = No. of moles / n-factor / Vol in litres.
So, 0.02 = n x 2/0.1 i.e. n = 0.001
Thus no. of molecules = Avagadro no / No. of moles = 6.022 x 1020 molecules